OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. Type your answer in the box without skipping any spaces. m = -4, point (1, -3)

OpenStudy (anonymous):

\[y-y_1=m(x-x_1) \rightarrow y-(-3)=-4(x-1) \rightarrow y+3=-4x-4 \rightarrow y=-4x-7\]